Your company uses the quadratic model y = -4.5x^2 + 150x to represent the average number of new customers who will be signed on (x) weeks after the release of your new service. How many new customers can you expect to gain in week 28?
This equation looks linear, but it says that it is quadratic, so I am not sure what do I start with. Thanks for your help on this.

Answer by nerdybill(7384)

Just plug x=28 into:
y = -4.5x^2 + 150x
.
y = -4.5(28)^2 + 150(28)
y = -4.5(784) + 4200
y = -3528 +4200
y = 672 customers